OpenStudy (anonymous):

If f(3)=9, f'(3)=-5, g(3)=3,g'(3)=-4, find F'(3) if F(x)=((f(x)g(x))^(1/3)

OpenStudy (anonymous):

chain rule plus product rule

OpenStudy (anonymous):

\[F'(x)=\frac{1}{3}\left(fg\right)^{-\frac{2}{3}}\times \left(f'g+g'f\right)\] plug in the numbers see what you get